# Real-Time Sampling-Based Safe Motion Planning for Robotic Manipulators in Dynamic Environments

### v1

@@ V Simulation Study
This section provides an extensive simulation study [^2] The implementation of DRGBT algorithm in C++ is available here. We use The Kinematics and Dynamics Library (KDL), Flexible Collision Library (FCL) [53] – optional, and nanoflann library [62]. The simulation was performed using the laptop PC with Intel®  Core™  i7-9750H CPU @ 2.60 GHz $\times$ 12 with 16 GB of RAM. with two distinct goals. The first goal is to subject DRGBT to a large number of scenarios to establish the appropriate deadline allocation for a real-time scheduling algorithm. Another goal is to compare DRGBT to a state-of-the-art method within a randomized trial.
@@ V-A Scenario Setup
1) Obstacles are placed w.r.t. uniform distribution within $\mathcal{W}$. For the sake of simplicity, we assume that each obstacle is an axis-aligned bounding box (AABB) with fixed dimensions of $1\times 1\times 1\,\mathrm{[cm^{3}]}$. Each obstacle is assigned a random constant velocity, where its magnitude is limited to [^3] We choose such value according to the ISO TS 15055 recommendations, which specify that in human-robot collaboration scenarios, if the human’s velocity is not being monitored, the system should assume a velocity of $1.6\,\mathrm{[\frac{m}{s}]}$ in the direction that minimizes the human-robot separation distance. $1.6\,\mathrm{[\frac{m}{s}]}$.
2) The used workspace $\mathcal{W}$ is bounded by a sphere with the radius of $1.5\,\mathrm{[m]}$ and the center in $(0,0,0.267)\,\mathrm{[m]}$, since this point represents the origin of the second robot’s joint frame (i.e., it is the robot’s base top point). If an obstacle reaches the limit of $\mathcal{W}$ or gets closer to $\max\{\frac{v_{obs}}{\omega_{max,1}},r_{1}\}$ from the robot’s base with a radius $r_{1}$, it just bounces back to $\mathcal{W}$ with a random direction maintaining the velocity magnitude $v_{obs}$. In other words, random obstacles cannot exist in close vicinity of the manipulator’s base to prevent the inevitable collisions.
3) The configurations $\boldsymbol{q}_{start}$ and $\boldsymbol{q}_{goal}$ are generated randomly in each run, with the condition that $\rho(\boldsymbol{q}_{start},\boldsymbol{q}_{goal})>\rho_{0}$, with $\rho$ being a $\mathcal{C}$-space metric from [4], and $\rho_{0}$ a proper threshold. Fig. 10 depicts an example of $\boldsymbol{q}_{start}$ alongside with 10 random obstacles (red boxes) and their velocity vectors illustrated in RViz (left) and Gazebo (right). The manipulator is mounted on the round green table with the radius of $67\,\mathrm{[cm]}$ as shown in Gazebo, which is also considered a static obstacle.
@@
  • the critical distance is $d_{crit}=0.05\,[\mathrm{m}]$;
@@
Figure 10: An example of the initial configuration alongside with 10 random obstacles in RViz (left) and Gazebo (right)
@@
Figure 11: Corresponding CDFs for the routines: $\mathtt{replan}$, $\mathtt{computeDistances}$, $\mathtt{generateGBur}$, $\mathtt{generateHorizon}$, $\mathtt{updateHorizon}$ and $\mathtt{updateCurrState}$. Maximal/execution times are indicated in each subfigure. Legend: DRGBT (red) and DRGBT-safe (green).
@@ V-B Simulation without Time Constraints
To simplify the time analysis of each task, we decompose it into routines that are analyzed separately. More precisely, each line from Alg. 1 represents a single routine (a part of a task), which is characterized by its execution/maximal time. The black lines indicate that the corresponding routine consumes negligible time (less than $1\,\mathrm{[\mu s]}$). The blue and red lines require substantial time and are subjected to the following analysis.
Fig. 11 shows corresponding cumulative density functions (CDF) for the routines $\mathtt{replan}$, $\mathtt{computeDistances}$, $\mathtt{generateGBur}$, $\mathtt{generateHorizon}$, $\mathtt{updateHorizon}$ and $\mathtt{updateCurrState}$. It reveals the completion success of each routine within the desired time. Maximal times are also indicated since graphics’ views are limited to the region of interest. Moreover, mean response time and standard deviation w.r.t. $N_{obs}$ for the same routines are provided by Fig. 12. The procedures $\mathtt{generateHorizon}$, $\mathtt{updateHorizon}$ and $\mathtt{updateCurrState}$ do not seem to be scenario-dependent (depicted by blue lines in Alg. 1) while other routines are (depicted by red lines in Alg. 1). The response time is approximately linear w.r.t. $N_{obs}$. Note that the abscissa scale is not linear since the considered numbers of obstacles in separate scenarios do not form the arithmetic sequence. It is worth stressing that Figs. 11 and 12 show data for DRGBT (-safe) in red (green). Since the only difference appears in the $\mathtt{updateCurrState}$ routine, it is clearly indicated.
@@
Generally, the worst-case scenario for the task occurs when its response time is equal to its execution time. In this regard, Fig. 11 reveals that maximal time of T1 is $e_{1}\approx 70\,\mathrm{[ms]}$ for DRGBT, and $e_{1}\approx 73\,\mathrm{[ms]}$ for DRGBT-safe, at least for the considered scenarios. In extreme cases involving more than 50 obstacles, $e_{1}$ will likely become larger. To summarize, the most time-consuming routines are $\mathtt{replan}$ and $\mathtt{generateGBur}$, which calls for their dedicated treatment in the sequel.
@@ V-C Imposing Time Constraints
Figure 13: Criteria w.r.t. $N_{obs}$ for different values of $T$ (shown in legend in $\mathrm{[ms]}$) in cases $u_{1}=1$ (top) and $u_{1}=0.5$ (bottom)
@@ V-D Simulation with Imposed Time Constraints
[…]
@@
Accounting for all criteria previously discussed, the convenient choice for the DRGBT(1) period is $T^{*}=50\,\mathrm{[ms]}$, implying the algorithm hard real-time frequency $f_{alg}^{*}=20\,\mathrm{[Hz]}$. This choice turns out as convenient for real-world applications as will be shown in Sec. VI. The choice of $f_{alg}^{*}$ is also tightly related to typical frame rates ( $\sim\,25\,\mathrm{[Hz]}$) of commercially available depth cameras. Unequivocally, the obtained results for DRGBT(2) suggest using $T^{*}=90\,\mathrm{[ms]}$ yielding a relatively low frequency of $f_{alg}^{*}=11.1\,\mathrm{[Hz]}$. Overall results from Fig. 14 suggest that using RGBMT* method in the replanning procedure would yield better performance than with RGBT-Connect, particularly in terms of the algorithm frequency.
Figure 15: Criteria w.r.t. $N_{obs}$ for DRGBT(1) and DRGBT(2) with and without replanning procedure (solid and dotted line, respectively)
@@
V-E Comparison to State-of-the-art Method
In [50], the preliminary version of DRGBT was shown to outperform RRTX [30] within a variety of scenarios. In this paper, we opt to compare the upgraded version of DRGBT to one of the recently proposed state-of-the-art algorithms. As the most representative example, we choose MARS algorithm [46], primarily due to its proven performance in comparison with other competing algorithms (such as DRRT [15], Anytime DRRT [20], and MPRRT [24]), and the availability of its C++ code implementation. The iteration time is limited to $50\,\mathrm{[ms]}$ (since $f_{alg}^{*}=20\,\mathrm{[Hz]}$). The scenario setup remains the same as described in Subsec. V-A, i.e., 15000 runs for each method.
[…]
Regarding algorithm time, DRGBT consistently finds solutions faster. However, the traversed path is relatively shorter for MARS, which can be attributed to the fact that it spends considerable time ($0.64\pm 0.64\,\mathrm{[s]}$) upfront to generate several initial optimal paths using RRT* method. The dashed lines for these two criteria indicate results restricted to those runs where both algorithms successfully find a solution. Clearly, both criteria are enhanced for DRGBT, while not for MARS, proving DRGBT’s effectiveness in relatively difficult scenarios.
[…]
Figure 16: Criteria for DRGBT ($D$) and MARS ($M$) w.r.t. $N_{obs}$; Legend left: —– $D$, —– $M$, - - - $D\backslash M$, - - - $M\backslash D$, - - - $D\cap M$, —– $D\cup M$; middle and right: —– $D$ and —– $M$ with their confidence intervals, - - - $D$ and - - - $M$ for runs where both methods succeed.
VI Experiments

### v2

@@ V Simulation Study
This section provides an extensive simulation study [^2] The implementation of DRGBT algorithm in C++ is available here. We use The Kinematics and Dynamics Library (KDL), Flexible Collision Library (FCL) [53], and nanoflann library [62]. It is worth mentioning that using FCL is optional since we implemented our library which is capable of fast handling simple primitive objects such as boxes, spheres and capsules. The simulation was performed using the laptop PC with Intel®  Core™  i7-9750H CPU @ 2.60 GHz $\times$ 12 with 16 GB of RAM, with the code compiled to run on a single core of the CPU without any GPU acceleration. with two distinct goals. The first goal is to subject DRGBT to a large number of scenarios to establish the appropriate deadline allocation for a real-time scheduling algorithm. Another goal is to utilize such scenarios to compare DRGBT to state-of-the-art methods within a randomized trial through Sec. VI.
@@ V-A Scenario Setup
Figure 10: An example of the initial configuration alongside 10 random obstacles in RViz (left) and Gazebo (right)
Figure 11: Corresponding CDFs for the routines: $\mathtt{replan}$, $\mathtt{computeDistances}$, $\mathtt{generateGBur}$, $\mathtt{generateHorizon}$, $\mathtt{updateHorizon}$ and $\mathtt{updateCurrState}$. Execution times are indicated in each subfigure. Legend: DRGBT (red) and DRGBT-safe (green).
1) Obstacles are placed w.r.t. uniform distribution within $\mathcal{W}$. For the sake of simplicity, we assume that each obstacle is an axis-aligned bounding box (AABB) with fixed dimensions of $1\times 1\times 1\,\mathrm{[cm^{3}]}$. Each obstacle is assigned a random constant velocity, where its magnitude is limited to [^3] We choose such value according to the ISO TS 15055 recommendations, which specify that in human-robot collaboration scenarios, if the human’s velocity is not being monitored, the system should assume a velocity of $1.6\,\mathrm{[\frac{m}{s}]}$ in the direction that minimizes the human-robot separation distance. $1.6\,\mathrm{[\frac{m}{s}]}$. The choice of these relatively small dimensions stems from the motivation to increase the likelihood that the algorithm finds a solution in case of many obstacles (e.g., 50). The primary objective of this study is to evaluate how the algorithm scales with the number of obstacles processed. A large obstacle count may also arise from decomposing a few relatively larger obstacles into multiple geometric primitives (e.g., AABBs). Apparently, larger obstacles cause frequent obstructions of goal configuration and thus decimating the success rates of tested algorithms. Conversely, scenarios with small obstacles are not necessarily trivial, especially when dealing with a large number of them. Since the robot occupies significant volume (particularly when approximated by bounding volumes, e.g., capsules), even fast moving point-like obstacles pose a substantial threat for collision.
2) The used workspace $\mathcal{W}$ is bounded by a sphere with the radius of $1.5\,\mathrm{[m]}$ and the center in $(0,0,0.267)\,\mathrm{[m]}$, since this point represents the robot’s base top point. If an obstacle reaches the limit of $\mathcal{W}$ or gets closer to $\max\{\frac{v_{obs}}{\omega_{max,1}},r_{1}\}$ from the robot’s base with a radius $r_{1}$, it just bounces back to $\mathcal{W}$ respecting the law of reflection while maintaining the velocity magnitude $v_{obs}$. In other words, random obstacles cannot exist in close vicinity of the manipulator’s base to prevent the inevitable collisions.
3) The configurations $\boldsymbol{q}_{start}$ and $\boldsymbol{q}_{goal}$ are generated randomly in each run, with the condition $\rho(\boldsymbol{q}_{start},\boldsymbol{q}_{goal})>\rho_{0}$, with $\rho$ being a $\mathcal{C}$-space metric from [4], and $\rho_{0}$ a proper threshold. Fig. 10 depicts an example of $\boldsymbol{q}_{start}$ alongside 10 random obstacles (red boxes) and their velocity vectors illustrated in RViz (left) and Gazebo (right). The manipulator is mounted on the round green table with the radius of $67\,\mathrm{[cm]}$ as shown in Gazebo, which is also considered a static obstacle.
@@
  • the critical workspace distance is $d_{crit}=0.05\,[\mathrm{m}]$;
@@ V-B Simulation without Time Constraints
To simplify the time analysis of each task, we decompose it into routines that are analyzed separately. More precisely, each line from Alg. 1 represents a single routine (a part of a task), which is characterized by its execution time. The black lines indicate that the corresponding routine consumes negligible time (less than $1\,\mathrm{[\mu s]}$). The blue and red lines require substantial time and are subjected to the following analysis.
Fig. 11 shows corresponding cumulative density functions (CDFs) for the routines $\mathtt{replan}$, $\mathtt{computeDistances}$, $\mathtt{generateGBur}$, $\mathtt{generateHorizon}$, $\mathtt{updateHorizon}$ and $\mathtt{updateCurrState}$. It reveals the completion success of each routine within the desired time. Maximum times are also indicated since graphics’ views are limited to the region of interest. Moreover, mean response time and standard deviation w.r.t. $N_{obs}$ for the same routines are provided by Fig. 12. The procedures $\mathtt{generateHorizon}$, $\mathtt{updateHorizon}$ and $\mathtt{updateCurrState}$ do not seem to be scenario-dependent (depicted by blue lines in Alg. 1) while other routines are (depicted by red lines in Alg. 1). The response time is approximately linear w.r.t. $N_{obs}$. Note that the abscissa scale is not linear since the considered numbers of obstacles in separate scenarios do not form the arithmetic sequence. It is worth stressing that Figs. 11 and 12 show data for DRGBT (-safe) in red (green). Since the only difference appears in the $\mathtt{updateCurrState}$ routine, it is clearly indicated.
@@
Generally, the worst-case scenario for the task occurs when its actual response time is equal to its execution/maximum time. In this regard, Fig. 11 reveals that maximum time of T1 is $e_{1}\approx 70\,\mathrm{[ms]}$ for DRGBT, and $e_{1}\approx 73\,\mathrm{[ms]}$ for DRGBT-safe, at least for the considered scenarios. In extreme cases involving more than 50 obstacles, $e_{1}$ will likely become larger. To summarize, the most time-consuming routines are $\mathtt{replan}$ and $\mathtt{generateGBur}$, which calls for their dedicated treatment.
@@ V-C Imposing Time Constraints
Figure 13: Criteria w.r.t. $N_{obs}$ for different values of $T$ (shown in legend in $\mathrm{[ms]}$) in cases $u_{1}=1$ (top) and $u_{1}=0.5$ (bottom).
@@ V-D Simulation with Imposed Time Constraints
Fig. 13 shows how decreasing $T$ (period of T1) from $120\,\mathrm{[ms]}$ towards $10\,\mathrm{[ms]}$ affects different criteria used for assessing the performance of DRGBT, such as success rate, algorithm time and path length (denoted as criteria in the sequel) w.r.t. $N_{obs}$. Three top figures provide results in case when the utility of T1 can go up to $u_{1}=1$, i.e., it may consume all the available time from the processor implying $e_{1}=T$. Similar results are obtained by three bottom figures, except the utility of T1 can go up to $u_{1}=0.5$ meaning $e_{1}=0.5T$. In both cases, T2 may be executed as needed during the remaining time after T1 completion respecting (3). Clearly, when $e_{1}<70\,\mathrm{[ms]}$ in Fig. 13, the scheduler may interrupt T1 execution, which makes it interesting to inspect potential consequences. It is worth mentioning that the success rate for zero random obstacles can be less than $100\,[\%]$, since the maximal available runtime of DRGBT can be exceeded.
Averaged results for criteria over all scenarios w.r.t. $T$ are shown in Fig. 14 for DRGBT(1) (using RGBMT* as replanner) and DRGBT(2) (using RGBT-Connect as replanner, where the same simulation study is carried out as for DRGBT(1)). Clearly, decreasing $T$ provides more timely reactions to the changing environment, which yields a higher success rate. On the other hand, we cannot decrease it to zero, since it would cause the decrease of $e_{1}$ at the same time, which may interrupt T1 more frequently. Thus, we identify the “sweet spot”, which lies around $80\,\mathrm{[ms]}$ for DRGBT(1), and around $90\,\mathrm{[ms]}$ for DRGBT(2). As for the averaged time and path length of DRGBT(1), “sweet spot” can be found around $30\,\mathrm{[ms]}$. CDFs from Fig. 11 reveal that $\mathtt{generateGBur}$ can be successfully completed in $99.9859\,[\%]$ cases within $30\,\mathrm{[ms]}$, while all other routines (except $\mathtt{replan}$) complete in $100\,[\%]$ cases. On the other hand, the best averaged time and path length of DRGBT(2) lies around $90\,\mathrm{[ms]}$.
@@
Accounting for all criteria previously discussed, the convenient choice for the DRGBT(1) period is $T^{*}=50\,\mathrm{[ms]}$, implying the algorithm hard real-time frequency $f_{alg}^{*}=20\,\mathrm{[Hz]}$. This choice turns out as convenient for real-world applications as will be shown in Sec. VII. The choice of $f_{alg}^{*}$ is also tightly related to typical frame rates ( $\sim\,25\,\mathrm{[Hz]}$) of commercially available depth cameras. Unequivocally, the obtained results for DRGBT(2) suggest using $T^{*}=90\,\mathrm{[ms]}$ yielding a relatively low frequency of $f_{alg}^{*}=11.1\,\mathrm{[Hz]}$. Overall results from Fig. 14 suggest that using RGBMT* method in the replanning procedure would yield better performance than with RGBT-Connect, particularly in terms of the algorithm frequency.
Figure 15: Criteria w.r.t. $N_{obs}$ for DRGBT(1) and DRGBT(2) with and without replanning procedure (solid and dotted line, respectively).
Averaged results for DRGBT-safe according to criteria w.r.t. $T$ are also depicted in Fig. 14. Clearly, the success rate is lower as expected, and decreases with increasing $T$. Nevertheless, it is worth reminding that all occurred collisions are type II collisions. In cca $\mathrm{2.94\,[\%]}$ cases, the maximal runtime is exceeded and those runs are labeled unsuccessful. Average algorithm time has significantly increased as expected since DRGBT-safe typically generates shorter trajectories/splines in order to ensure their existence within dynamic expanded bubbles. Consequently, a price to pay is a slower motion of the manipulator which may trigger collisions more frequently. Interestingly, average path length has improved, since adhering to the predefined path is usually the safest motion, which means the robot possibly does not explore potentially unsafe regions. To summarize, according to all criteria, DRGBT-safe performs better for smaller values of $T$ ($10$ to $20\,\mathrm{[ms]}$) implying higher algorithm frequencies $f_{alg}^{*}\in[50,100]\,\mathrm{[Hz]}$).
VI Comparison to State-of-the-art Methods
In [50], the preliminary version of DRGBT was shown to outperform RRTX algorithm [30] within a variety of scenarios. For the sake of completeness, we include RRTX in the comparative study [^4] To this end, we have implemented a C++ version of the algorithm available here, by carefully following the original Julia code made available by the author here. Since the code is entirely reimplemented, we equipped it with the same trajectory parametrization as DRGBT in order to make sure that all kinematic constraints are satisfied.. Moreover, we compare DRGBT to one of the recently proposed state-of-the-art algorithms. As the most representative example, we choose MARS algorithm [46], primarily due to its proven performance in comparison with other competing algorithms (such as DRRT [15], Anytime DRRT [20], and MPRRT [24]), and the availability of its C++ implementation. Each method uses its own default parameters.
VI-A Randomized trial
@@
Regarding algorithm time, DRGBT consistently finds solutions faster. However, the traversed path is approximately the same for DRGBT and MARS, while RRTX yields significantly longer paths. Clearly, DRGBT outperforms the competing algorithms w.r.t. both criteria, proving its effectiveness in relatively difficult scenarios.
Overall, DRGBT demonstrates superior performance. Please note that, unlike DRGBT, MARS relies on the following:
  • Replanning is executed in parallel (at the frequency of $5\,\mathrm{[Hz]}$) with all other tasks of the algorithm;
  • While searching for the initial paths, obstacles remain stationary. The algorithm consumes considerable time ($0.53\pm 0.68\,\mathrm{[s]}$) upfront to generate several initial paths;
  • Kinematic constraints $\mathcal{K}$ are not explicitly considered during trajectory design, leading to violations of the maximal joint velocity in $10.5\,[\%]$ runs (with the values of $3.45\pm 0.22\,\mathrm{[\frac{rad}{s}]}$), and the maximal joint acceleration in $34.2\,[\%]$ runs (with the values of $177.9\pm 140.5\,\mathrm{[\frac{rad}{s^{2}}]}$).
In this context, the success rate of MARS can be regarded as “conditional”. We presume that MARS algorithm can be modified to account for kinematic constraints. It remains an open question whether such modification would result in performance deterioration. Nevertheless, we refrained from intervening on the original code. On the other hand, if we relax the acceleration and jerk constraints for DRGBT, the obtained results become substantially better (e.g., success rate increases from $25\,\mathrm{[\%]}$ to $60\,\mathrm{[\%]}$ for $N_{obs}=50$).
Figure 16: Criteria for DRGBT ($D$) and MARS ($X=M$) – top, and RRTX ($X=R$) – bottom, w.r.t. $N_{obs}$.
VI-B Scenarios with larger obstacles and narrow passages
In order to better evaluate the algorithm’s performance, two more complex scenarios including larger obstacles and narrow passages, as shown in Fig. 17, are added within the simulation comparison study. The scenarios are based on those from [50], this time set up consistently with the real-time implementation. The robot workspace is shown, where the initial and goal configuration are indicated (magenta and green, respectively). The obstacles (black) and their paths are also clearly indicated (gray with red arrows). In the first scenario, four prismatic obstacles follow a circular path in a horizontal plane. To reach the goal, the robot needs to squeeze through two moving adjacent obstacles. In the second scenario, the obstacles are two deformable tunnels. The first tunnel is periodically stretching and shrinking, while the second one is moving left and right along the $x$-axis. The robot end-effector needs to escape one tunnel and squeeze into another one. In both scenarios, obstacles’ linear velocity is picked randomly from the range $(0,0.3]\,\mathrm{[\frac{m}{s}]}$ for each of 1000 algorithm’s runs. The maximal runtime is limited to $10\,\mathrm{[s]}$ for Scenario 1 and to $30\,\mathrm{[s]}$ for Scenario 2.
Tab. II shows the obtained results from the carried out study, where the success rate $\eta$, mean algorithm time $T_{alg}$, and mean path length $D$ are considered. Clearly, DRGBT outperforms both methods according to success rate, despite MARS violating kinematic constraints (e.g., on maximal joint velocity in $81.4\,\mathrm{[\%]}$ and $67.8\,\mathrm{[\%]}$ runs in scenarios 1 and 2, respectively; and on maximal joint acceleration and jerk in $100\,\mathrm{[\%]}$ runs in both scenarios). Scenario 2 is particularly difficult since the robot requires significant amount of time to escape one moving tunnel and to enter into another one. Tab. II proves that a trade-off between $\eta$, and $T_{alg}$ and $D$ exists. It is worth indicating that success rates of RRTX are now worse than those in [50] likely due to limited runtime.
Figure 17: Three additional scenarios used in the comparison study.
TABLE II: Results for Scenarios 1 and 2 from the comparison study.
Scenario 1 Scenario 2
DRGBT MARS RRTX DRGBT MARS RRTX
$\eta\,[\%]$ $\boldsymbol{94.5}$ $25.8$ 99 $\boldsymbol{51.6}$ $14$ $17.9$
$T_{alg}\,\mathrm{[s]}$ $\boldsymbol{6.5}\pm\boldsymbol{4.5}$ $7.2\pm 1.6$ $7.4\pm 3.4$ $16.5\pm 7.7$ $\boldsymbol{8.9}\pm\boldsymbol{1.9}$ $9.5\pm 4.4$
$D\,\mathrm{[rad]}$ $16.1\pm 11.8$ $\boldsymbol{13.1}\pm\boldsymbol{3.9}$ $13.5\pm 3.1$ $37.4\pm 16.8$ $\boldsymbol{14.4}\pm\boldsymbol{4.4}$ $18.4\pm 4.4$
VI-C Scalability beyond 6-DoFs
To investigate the impact of dimensionality on algorithm performance, the competing methods are evaluated on a spatial $n$-DoF manipulator, where $n\in\{10,18\}$ (Scenario 3 in Fig. 17 depicts spatial 14-DoF robot). The environment is populated with five randomly placed obstacles, each with fixed dimensions of $30\times 30\times 30\,\mathrm{[cm^{3}]}$. Obstacles motion undergoes the same principles as described in Subsec. V-A. For each of 1000 trials, the linear velocity of each obstacle is randomly chosen from the range $(0,0.3]\,\mathrm{[\frac{m}{s}]}$. The maximal allowable runtime per trial is fixed to $30\,\mathrm{[s]}$.
Tab. III shows the obtained results. In parentheses next to success rate, the percentage of runs is depicted when the algorithm runtime exceeded without colliding with obstacles. Such measure reveals that collisions for DRGBT occur in less than $4\,[\%]$ cases. Clearly, the proposed approach displays dominant performance in the case of 10-DoF robot. As for 18-DoF robot, DRGBT achieves a higher success rate than competing methods, but both $\eta$ and $T_{alg}$ turn out to be worse, which indicates a substantial decrease in these performance metrics w.r.t. number of joints $n$. Moreover, the practicality of all the algorithms become questionable since waiting for several minutes for the robot to reach the goal might be unacceptable. Regardless, DRGBT appears to be superior in term of success rate and avoiding collisions within the tested scenarios with higher values of $n$. To better support this claim, clearly a more comprehensive investigation is necessary over a wider range of manipulator types and obstacle counts.
TABLE III: Results for Scenario 3 from the comparison study.
10-DoF robot 18-DoF robot
DRGBT MARS RRTX DRGBT MARS RRTX
$\eta\,[\%]$ $\boldsymbol{89.1}$ ($\boldsymbol{7.3}$) $68$ ($2.1$) $44.6$ ($4.3$) $\boldsymbol{41.2}$ ($\boldsymbol{57.1}$) $37.5$ ($45.2$) $29.4$ ($48$)
$T_{alg}\,\mathrm{[s]}$ $\boldsymbol{8.6}\pm\boldsymbol{7.6}$ $12.1\pm 7.4$ $9.7\pm 4.9$ $16.9\pm 11.4$ $20.6\pm 7.7$ $\boldsymbol{10.1\pm 5.3}$
$D\,\mathrm{[rad]}$ $\boldsymbol{20.1}\pm\boldsymbol{14.4}$ $23.0\pm 12.4$ $21.2\pm 6.1$ $48.3\pm 29.3$ $41.9\pm 14.6$ $\boldsymbol{30.2\pm 7.4}$
Redundancy constitutes a critical factor in robotics, necessitating careful consideration. Many high-DoF manipulators exhibit kinematic redundancy, which can be exploited to optimize secondary objectives such as manipulability, energy efficiency, joint limit avoidance, and collision avoidance. Integrating redundancy resolution techniques, including null-space projection methods [63, 64] or optimization-based approaches [65], could further extend the applicability of the proposed DRGBT algorithm to redundant systems. The results obtained in relatively high-dimensional $\mathcal{C}$-spaces within Tab. III demonstrate the potential of DRGBT for effectively addressing redundancy. However, the present work exclusively formulates goal configurations within the robot’s $\mathcal{C}$-space. The incorporation of methods that define goals in the robot’s workspace is left for future investigation.
VII Experiments